The xy-plane intersects the sphere x^2+y^2+z^2+2x+2y-z=2 in a great circle.
True or false with full explanation
1
Expert's answer
2020-11-24T05:32:17-0500

The xy-plane intersects the sphere "x^2+y^2+z^2+2x+2y-z=2" in a great circle if and only if the center of this sphere belong to "xy"-plane. Let us rewrite the equation of the sphere in the the following form: "(x+1)^2+(y+1)^2+(z-\frac{1}{2})^2=2+1+1+\frac{1}{4}=\frac{17}{4}." It follows that "M(-1,-1,\frac{1}{2})" is the center of the sphere. Taking into account that the third coordinate of "M" is not equal to 0, we conclude that the center of the sphere does not belong to the "xy"-plane, and therefore, the sphere "x^2+y^2+z^2+2x+2y-z=2" does not intersect the "xy"-plane in a great circle.


Answer: false